i need help solving an equation with fractions in it. I have worked it out but i want to be sure my answer is correct.
Original problem 3-2/5x=6
i did -2/5x=6-3
-2/5x=3
x=3/1(divided)-5/2
x=3/1*-5/2=-15/2

Answer by Alan3354(69443)

Original problem 3-2/5x=6
i did -2/5x=6-3
-2/5x=3
x=3/1(divided)-5/2
x=3/1*-5/2=-15/2
------------
It's correct, but I would eliminate fractions first.
3 - (2/5)x = 6
Times 5:
15 - 2x = 30
-2x = 15
x = -15/2